a minor league team plays 113 games and won 17 more than 3 times as many games as they lost. what are the wins and losses for the season?
Answer by macston(5194)

a minor league team plays 113 games and won 17 more than 3 times as many games as they lost. what are the wins and losses for the season?
W=won; L=lost
W+L=113
W=113-L
W=3L+17 Substitute for W
113-L=3L+17 Subtract 17 from each side.
96-L=3L Add L to each side
96=4L Divide each side by L
24=L ANSWER 1: The team lost 24 games.
W=113-L=113-24=89 ANSWER 2: The team won 89 games
CHECK:
W=3L+17
89=3(24)+17
89=72+17
89=89